# Multi-Personality Generation of LLMs at Decoding-time

### v1

@@ 1. Introduction
In recent years, Large Language Models (LLMs) (Zhao et al., 2023) have witnessed rapid development, with parameter scales surging from billions to trillions and training data covering diverse domains such as text and code (Chen et al., 2021; Li et al., 2024), thereby endowing them with capabilities in reasoning and semantic understanding (Huang and Chang, 2022). Within this context, multi-personality generation has emerged as a critical task, which requires LLMs to transform from general-purpose tools into personalized intelligences capable of understanding and addressing the nuanced needs of individual users, and is widely applied in scenarios such as personalized question answering, artificial assistants, and intelligent customer service (Li et al., 2024; Guo et al., 2024; Cheng et al., 2024; Jia et al., 2024).
@@
Existing studies can be broadly categorized into retraining-based and decoding-time methods. Retraining methods seek to encode multiple preference dimensions into a single model during training via multi-objective optimization. This involves techniques like multi-objective reinforcement learning where models learn to balance competing preferences through iterative feedback and weighted-loss supervised fine-tuning, which assigns differential weights to preference signals to guide parameter updates (Harland et al., 2024; Zhou et al., 2023). But these approaches suffer from notable drawbacks: they incur high training costs and exhibit poor scalability when adapting to new preferences, often requiring full retraining. In contrast, decoding-time methods (Huang et al., 2024; Khanov et al., 2024; Sun et al., 2024) circumvent the need for retraining by guiding the decoding process through external tools. Specifically, they require a model capable of evaluating the quality of responses at the multi-personality level, and leverage this model as reward model  (Chen et al., 2024; Yang et al., 2024b; Khanov et al., 2024) or aligner  (Yang et al., 2024a) to shape output, or introduce target preference signals directly into generation via prompt learning (Chen et al., 2024; Ding et al., 2021). Yet, they depend heavily on external models that are hard to acquire and struggle to handle preferences that change dynamically.
To provide efficiency and practicality in dynamic scenarios where user preferences evolve, combining multiple models at decode time rather than relying on external reward models has emerged as a trend. Specifically, such approaches require the existence of a model for each dimension (e.g., the model for Hobby is a single-dimensional model in Figure 1). And they involve either combining model parameters across multi-dimensions  (Jang et al., 2023; Ramé et al., 2023; Lu et al., 2023) or linearly combining the prediction logits of models from multiple dimensions to generate the next token  (Shi et al., 2024). However, their combinatorial mechanisms remain heuristic, with performance constrained by the capabilities of individual constituent models and lacking robustness. Overall, there remains a lack of a model capable of flexibly integrating diverse personality dimensions while maintaining efficiency, scalability, and robust performance across dynamic scenarios.
@@ 2. Related Work
Reinforcement Learning from Human Feedback. RLHF (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a; Dubois et al., 2023; Bai et al., 2022b) serves as the dominant paradigm for aligning Large Language Models with human preferences. It traditionally employs a three-stage pipeline: Supervised Fine-Tuning (SFT), training a reward model from human feedback, and then enhancing the policy through a reinforcement learning algorithm, most notably PPO (Schulman et al., 2017). More recent approaches, such as DPO (Rafailov et al., 2023), offer an alternative to this final stage by directly optimizing the policy on preference data, bypassing the need for an explicit reward model and complex RL optimization.
Decoding-time Alignment. Decoding-time alignment methods emerged as an effective approach to satisfy diverse user requirements which can adapt to different objectives without retraining. PAD (Chen et al., 2024) introduces token-level personalized rewards by decoupling text generation dynamics from user preferences, enabling dynamic adjustment of base model predictions during inference. This approach leverages a single policy and reward model to generalize to unseen preferences, outperforming training-based methods in multi-objective alignment while remaining computationally efficient. MOD (Shi et al., 2024) employs a closed-form solution derived via Legendre transformation to combine predictions from multiple base models-each aligned to distinct objectives-during decoding. By weighting model outputs algebraically, MOD achieves precise control over arbitrary objective weightings, demonstrating scalability across model sizes and robustness to unseen preference combinations.
Multi-dimensional Personalization. For the more complex multi-dimensional personalization alignment challenge of having models satisfy multiple (or even conflicting) objectives simultaneously (e.g., balancing usefulness and harmlessness, or simulating a specific MBTI/Role-Playing persona), the research community has explored different strategies. The main approaches include training a single model to optimize a weighted multi-objective function (e.g., MORLHF, MODPO (Zhou et al., 2023)), combining the parameters of independently-trained single-preference models (e.g., DPO Soups (Jang et al., 2023)) during decoding, and bootstrapping by combining reward signals or model predictions at decoding (e.g., PAD (Chen et al., 2024), MOD (Shi et al., 2024)).
@@ 3.1. Problem Definition and Formulation
For $\pi_{d_{i}}(y|x)$, the RLHF (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a; Dubois et al., 2023) training objective under $f$-divergence regularization for the reward functions $\mathcal{R}_{i}$, the closed form of $\pi_{d_{i}}$ can be written as:
@@
where $\beta>0$ is a regularization parameter. For the preference on dimension $i$, our single-attribute preference model $\pi_{d_{i}}(y|x)$ can be viewed as an optimal policy trained to maximize an implicit reward function ${\mathcal{R}}_{i}(y|x)$. As stated in the work of MOD (Shi et al., 2024), if the barrier function $f$ is continuously differentiable and strongly convex on $\mathbb{R}_{+}$, the following closed-form bijection between $\pi_{d_{i}}$ and the corresponding $\mathcal{R}_{i}$ can be obtained:
@@
Unlike approaches that directly optimize the combined reward during training time via reinforcement learning (Ziegler et al., 2019; Shi et al., 2024; Ouyang et al., 2022; Bai et al., 2022a; Dubois et al., 2023; Kim et al., 2023), our method draws inspiration from the mixture of models during decoding (Wortsman et al., 2022; Lu et al., 2023; Benabbou and Perny, 2015). We directly map the optimization objective of the combined reward to the aggregation of various preference models at the policy level, and define the multi-personality generation objective as sampling from a target probability distribution $\pi_{\mathrm{MPG}}(y|x)$.
@@ 3.2. MPG Based on Density Ratio
The core principle guiding our approach stems from a key feature observed in LLM preference alignment training (e.g., DPO (Rafailov et al., 2023)): the probability density ratio of an optimized policy model $\pi_{\theta}(y|x)$ relative to the reference model $\pi_{\text{ref}}(y|x)$ used in its training, $r(y|x)=\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$, implicitly encodes the preference information learned by the model. This ratio quantifies the model’s preference for a particular output sequence $y$ compared to its likelihood under $\pi_{\text{ref}}(y|x)$. We take advantage of this principle to sample from our target multi-personality distribution $\pi_{\mathrm{MPG}}(y|x)$.
@@
Although it is difficult to calculate $Z(x)$, the core idea of MPG is to apply rejection sampling to avoid these complex calculations. Its acceptance probability $A(y|x)$ is proportional to the ratio of the target distribution to the proposal distribution, that is:
(8) $A(y|x)=\frac{\pi_{\mathrm{MPG}}(y|x)}{M\cdot\pi_{\text{ref}}(y|x)}\propto\frac{\pi_{\mathrm{MPG}}(y|x)}{\pi_{\text{ref}}(y|x)}$
Combine Eq (7), for a giver inputxx, $Z(x)$ is a constant that does not change with $y$. Therefore, in the expression above of the target density ratio, $1/Z(x)$, as a multiplicative factor that is the same for all candidates $y$, its effect can be absorbed into the upper bound $M$ of the rejection sampling, the $1/\beta$ is also. And since $(\nabla f)^{(-1)}$ is usually monotonically increasing, it does not change the relative order of its arguments, so that we can get:
@@
Defining the individual attribute density ratio as $r_{i}(y|x)=\frac{\pi_{d_{i}}(y|x)}{\pi_{\text{ref}}(y|x)}$, representing the preference signal for the $i$-th attribute relative to the reference model. In most policy optimizations (the regularization terms are PPO (Schulman et al., 2017) and DPO (Rafailov et al., 2023)), the Reverse KL-divergence is used, with $f(x)=x\log x$ and $\nabla f(x)=\log x+1$. Therefore, we follow this to derive the formula for the density ratio of the target distribution relative to $\pi_{\text{ref}(y|x)}$ as follows:
@@
(11) $A(y|x)\propto\frac{\pi_{\mathrm{MPG}}(y|x)}{\pi_{\text{ref}}(y|x)}\propto\sum_{i=1}^{N}\alpha_{i}{r}_{i}(y|x)$
This relationship forms the theoretical cornerstone of our MPG method. It demonstrates that within our rejection sampling framework, the acceptance probability $A(y|x)$ for a given text sequence $y$ is proportional to the weighted sum of the base density ratios, $r_{i}(y|x)$, from each attribute preference model. This efficient representation explicitly encodes multi-dimensional preference information, enabling effective control over personality generation.
@@ 3.3. Speculative Chunk-level based Rejection sampling (SCR)
We implement MPG in a decoding-time algorithm termed Speculative Chunk-level based Rejection sampling (SCR). Unlike sequence-level rejection sampling, SCR proposes and validates chunks of $k$ tokens at a time, integrating speculative decoding (Leviathan et al., 2022) with chunk-wise acceptance decisions. By validating multi-token speculative proposals against the aggregated density ratio, SCR preserves the correctness of rejection sampling while amortizing expensive target-policy validation over longer spans, thereby reducing the frequency of preference model evaluations and enabling efficient parallelization.
@@ 4.1.1. MBTI Personality Simulation
For the MBTI Personality Simulation task, on the training stage, we construct a specialized training dataset derived from pandalla/Machine_Mindset_MBTI_dataset  [^2] https://huggingface.co/datasets/pandalla/Machine_Mindset_MBTI_dataset that captures the four fundamental MBTI dimensions $D_{\text{MBTI}}$ which contains data pairs $(y_{w},y_{l})_{i}$, where $y_{w}$ is preferred over $y_{l}$ on the $i$-th dimension. For evaluation, we employ three curated benchmarks: 1) MBTI-QA  [^3] https://huggingface.co/datasets/pandalla/Machine_Mindset_MBTI_dataset for instruction-following question answering, 2) MBTI-MCQA (Pan and Zeng, 2023) for multiple-choice answering, and 3) MBTI-16P  [^4] https://www.16personalities.com/free-personality-test containing items from the 16Personalities psychometric instrument. These datasets systematically assess personality-specific response patterns through carefully constructed diagnostic questions.
@@ 4.1.2. Role-Playing
For the Simple Role-Playing task, we constructed the dataset by adapting the methodology from ALOE  [^5] https://github.com/ShujinWu-0814/ALOE, focusing on user profile and target personality. The training set includes annotations generated by the ChatGPT-4o model (Hurst et al., 2024). The evaluation set, a subset of this dataset, contains conversation contexts grounded in user profile descriptions and specific persona. Generated responses were evaluated based on their ability to effectively simulate the described profile and persona within these contexts.
@@
We adopt the same evaluation methodology as for the MBTI personality simulation task, using LLM-as-a-Judge (Zheng et al., 2023). The evaluation dimensions are as follows:
@@ 4.2. Baselines
Preference Prompting (PP) (Jang et al., 2023) implements personality conditioning through explicit attribute descriptions in prompt engineering, applied across base model, single-preference model, and specialized model(use in Sec. 4.6) configurations. DPO Soups (a variant of Personalized Soups) (Jang et al., 2023; Zhou et al., 2023) acquires the LoRA parameters of independently - trained single - preference DPO models. Following weighted averaging, these parameters are applied to $\pi_{\text{ref}}$ to generate text with multi-personality. MOD (Shi et al., 2024) performs a weighted linear combination of the output logits from single - attribute DPO models during decoding and samples based on the combined logits. SCR: Our Speculative Chunk-level based Rejection sampling with parallel multi-preference scoring.
@@ 4.6. Leveraging Specialized Reference Models
To investigate this, we evaluated the performance of SCR when using a different, more specialized reference model. Specifically, we replaced the default base model with a strong, open-source model pre-trained for Role-Playing (Cui et al., 2023; Wang et al., 2025) [^6] https://modelscope.cn/organization/FarReelAILab, [^7] https://huggingface.co/Neph0s/CoSER-Llama-3.1-8B, while keeping our DPO-trained preference models $\pi_{d_{i}}$ unchanged.

### v2

@@ 1. Introduction
In recent years, Large Language Models (LLMs) (Zhao2023ASO) have witnessed rapid development, with parameter scales surging from billions to trillions and training data covering diverse domains such as text and code (Chen2021EvaluatingLL; Li2024PersonalLA), thereby endowing them with capabilities in reasoning and semantic understanding (Huang2022TowardsRI). Within this context, multi-personality generation has emerged as a critical task, which requires LLMs to transform from general-purpose tools into personalized intelligences capable of understanding and addressing the nuanced needs of individual users, and is widely applied in scenarios such as personalized question answering, artificial assistants, and intelligent customer service (Li2024PersonalLA; Guo2024LargeLM; Cheng2024ExploringLL; NEURIPS2024_1cb57fcf).
@@
Existing studies can be broadly categorized into retraining-based and decoding-time methods. Retraining methods seek to encode multiple preference dimensions into a single model during training via multi-objective optimization. This involves techniques like multi-objective reinforcement learning where models learn to balance competing preferences through iterative feedback and weighted-loss supervised fine-tuning, which assigns differential weights to preference signals to guide parameter updates (Harland2024AdaptiveAD; Zhou2023BeyondOA). But these approaches suffer from notable drawbacks: they incur high training costs and exhibit poor scalability when adapting to new preferences, often requiring full retraining. In contrast, decoding-time methods (Huang2024DeALDA; Khanov2024ARGSAA; NEURIPS2024_3950f6bf) circumvent the need for retraining by guiding the decoding process through external tools. Specifically, they require a model capable of evaluating the quality of responses at the multi-personality level, and leverage this model as reward model  (Chen2024PADPA; Yang2024RewardsinContextMA; Khanov2024ARGSAA) or aligner  (Yang2024MetaalignerTG) to shape output, or introduce target preference signals directly into generation via prompt learning (Chen2024PADPA; Ding2021OpenPromptAO). Yet, they depend heavily on external models that are hard to acquire and struggle to handle preferences that change dynamically.
To provide efficiency and practicality in dynamic scenarios where user preferences evolve, combining multiple models at decode time rather than relying on external reward models has emerged as a trend. Specifically, such approaches require the existence of a model for each dimension (e.g., the model for Hobby is a single-dimensional model in Figure 1). And they involve either combining model parameters across multi-dimensions  (Jang2023PersonalizedSP; Ram2023RewardedST; Lu2023InferringPF) or linearly combining the prediction logits of models from multiple dimensions to generate the next token  (Shi2024DecodingTimeLM). However, their combinatorial mechanisms remain heuristic, with performance constrained by the capabilities of individual constituent models and lacking robustness. Overall, there remains a lack of a model capable of flexibly integrating diverse personality dimensions while maintaining efficiency, scalability, and robust performance across dynamic scenarios.
@@ 2. Related Work
Reinforcement Learning from Human Feedback. RLHF (Ziegler2019FineTuningLM; Ouyang2022TrainingLM; Bai2022TrainingAH; Dubois2023AlpacaFarmAS; Bai2022ConstitutionalAH) serves as the dominant paradigm for aligning Large Language Models with human preferences. It traditionally employs a three-stage pipeline: Supervised Fine-Tuning (SFT), training a reward model from human feedback, and then enhancing the policy through a reinforcement learning algorithm, most notably PPO (Schulman2017ProximalPO). More recent approaches, such as DPO (Rafailov2023DirectPO), offer an alternative to this final stage by directly optimizing the policy on preference data, bypassing the need for an explicit reward model and RL optimization.
Decoding-time Alignment. Decoding-time alignment methods emerged as an effective approach to satisfy diverse user requirements, which can adapt to different objectives without retraining. PAD (Chen2024PADPA) introduces token-level personalized rewards by decoupling text generation dynamics from user preferences, enabling dynamic adjustment of base model predictions during inference. It leverages a single policy and reward model to generalize to unseen preferences. MOD (Shi2024DecodingTimeLM) employs a closed-form solution derived via Legendre transformation to combine predictions from multiple base models-each aligned to distinct objectives-during decoding. By weighting model outputs algebraically, MOD achieves precise control over objective weights. Drift (kim2025drift) extends this direction by inferring user preferences through attribute-level decomposition and combining attribute signals directly at decoding time, enabling effective few-shot personalization without explicit reward models.
Multi-dimensional Personalization. For the more complex multi-dimensional personalization alignment challenge of having models satisfy multiple (or even conflicting) objectives simultaneously (e.g., balancing usefulness and harmlessness, or simulating a specific MBTI/Role-Playing persona), the research community has explored different strategies. The main approaches include training a single model to optimize a weighted multi-objective function (e.g., MORLHF, MODPO (Zhou2023BeyondOA)), combining the parameters of independently-trained single-preference models (e.g., DPO Soups (Jang2023PersonalizedSP)) during decoding, and bootstrapping by combining reward signals or model predictions at decoding (e.g., PAD (Chen2024PADPA), MOD (Shi2024DecodingTimeLM)). Drift (kim2025drift) decomposes user preferences into interpretable attributes and composes them at decoding time, providing a new perspective on multi-dimensional personalization.
@@ 3.1. Problem Definition and Formulation
For $\pi_{d_{i}}(y|x)$, the RLHF (Ziegler2019FineTuningLM; Ouyang2022TrainingLM; Bai2022TrainingAH; Dubois2023AlpacaFarmAS) training objective under $f$-divergence regularization for the reward functions $\mathcal{R}_{i}$, the closed form of $\pi_{d_{i}}$ can be written as:
@@
where $\beta>0$ is a regularization parameter. For the preference on dimension $i$, our single-attribute preference model $\pi_{d_{i}}(y|x)$ can be viewed as an optimal policy trained to maximize an implicit reward function ${\mathcal{R}}_{i}(y|x)$. As stated in the work of MOD (Shi2024DecodingTimeLM), if the barrier function $f$ is continuously differentiable and strongly convex on $\mathbb{R}_{+}$, the following closed-form bijection between $\pi_{d_{i}}$ and the corresponding $\mathcal{R}_{i}$ can be obtained:
@@
Unlike approaches that directly optimize the combined reward during training time via reinforcement learning (Ziegler2019FineTuningLM; Shi2024DecodingTimeLM; Ouyang2022TrainingLM; Bai2022TrainingAH; Dubois2023AlpacaFarmAS; Kim2023AligningLL), our method draws inspiration from the mixture of models during decoding (Wortsman2022ModelSA; Lu2023InferringPF; 2015Incremental). We directly map the optimization objective of the combined reward to the aggregation of various preference models at the policy level, and define the multi-personality generation objective as sampling from a target probability distribution $\pi_{\mathrm{MPG}}(y|x)$.
@@ 3.2. MPG Based on Density Ratio
The core principle guiding our approach stems from a key feature observed in LLM preference alignment training (e.g., DPO (Rafailov2023DirectPO)): the probability density ratio of an optimized policy model $\pi_{\theta}(y|x)$ relative to the reference model $\pi_{\text{ref}}(y|x)$ used in its training, $r(y|x)=\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$, implicitly encodes the preference information learned by the model. This ratio quantifies the model’s preference for a particular output sequence $y$ compared to its likelihood under $\pi_{\text{ref}}(y|x)$. We take advantage of this principle to sample from our target multi-personality distribution $\pi_{\mathrm{MPG}}(y|x)$.
@@
Although it is difficult to calculate $Z(x)$, the core idea of MPG is to apply rejection sampling to avoid these complex calculations. The acceptance probability $A(y|x)$ is defined as the ratio between the target distribution and the proposal distribution scaled by an envelope constant $M$:
(8) $A(y\mid x)=\frac{\pi_{\mathrm{MPG}}(y\mid x)}{M\cdot\pi_{\mathrm{ref}}(y\mid x)}$
Where $M\geq 1$ is chosen such that $M\cdot\pi_{\mathrm{ref}}(y\mid x)\geq\pi_{\mathrm{MPG}}(y\mid x)$ holds for all $y$, ensuring that the acceptance probability always lies within $[0,1]$. Combine Eq (7), for a giver inputxx, $Z(x)$ is a constant that does not change with $y$. Therefore, in the expression above of the target density ratio, $1/Z(x)$, as a multiplicative factor that is the same for all candidates $y$, its effect can be absorbed into the upper bound $M$ of the rejection sampling, the $1/\beta$ is also. And since $(\nabla f)^{(-1)}$ is usually monotonically increasing, it does not change the relative order of its arguments, so that we can get:
@@
Defining the individual attribute density ratio as $r_{i}(y|x)=\frac{\pi_{d_{i}}(y|x)}{\pi_{\text{ref}}(y|x)}$, representing the preference signal for the $i$-th attribute relative to the reference model. In most policy optimizations (the regularization terms are PPO (Schulman2017ProximalPO) and DPO (Rafailov2023DirectPO)), the Reverse KL-divergence is used, with $f(x)=x\log x$ and $\nabla f(x)=\log x+1$. Therefore, we follow this to derive the formula for the density ratio of the target distribution relative to $\pi_{\text{ref}(y|x)}$ as follows:
@@
(11) $A(y|x)=\frac{\pi_{\mathrm{MPG}}(y|x)}{M\cdot\pi_{\text{ref}}(y|x)}\propto\frac{\sum_{i=1}^{N}\alpha_{i}{r}_{i}(y|x)}{M}$
This relationship forms the theoretical cornerstone of our MPG method. It demonstrates that within our rejection sampling framework, the acceptance probability $A(y\mid x)$ for a candidate text sequence is directly governed by the weighted sum of the base density ratios $r_{i}(y\mid x)$ contributed by the attribute-specific preference models. This efficient representation explicitly encodes multi-dimensional preference information, enabling effective control over personality generation.
@@ 3.3. Speculative Chunk-level based Rejection sampling (SCR)
We implement MPG in a decoding-time algorithm termed Speculative Chunk-level based Rejection sampling (SCR). Unlike sequence-level rejection sampling, SCR proposes and validates chunks of $k$ tokens at a time, integrating speculative decoding (Leviathan2022FastIF) with chunk-wise acceptance decisions. By validating multi-token speculative proposals against the aggregated density ratio, SCR preserves the correctness of rejection sampling while amortizing expensive target-policy validation over longer spans, thereby reducing the frequency of preference model evaluations and enabling efficient parallelization.
@@ 4.1.1. MBTI Personality Simulation
For the MBTI Personality Simulation task, on the training stage, we construct a specialized training dataset derived from pandalla/Machine_Mindset_MBTI_dataset  [^2] https://huggingface.co/datasets/pandalla/Machine_Mindset_MBTI_dataset that captures the four fundamental MBTI dimensions $D_{\text{MBTI}}$ which contains data pairs $(y_{w},y_{l})_{i}$, where $y_{w}$ is preferred over $y_{l}$ on the $i$-th dimension. For evaluation, we employ three curated benchmarks: 1) MBTI-QA  [^3] https://huggingface.co/datasets/pandalla/Machine_Mindset_MBTI_dataset for instruction-following question answering, 2) MBTI-MCQA (Pan2023DoLP) for multiple-choice answering, and 3) MBTI-16P  [^4] https://www.16personalities.com/free-personality-test containing items from the 16Personalities psychometric instrument. These datasets systematically assess personality-specific response patterns through carefully constructed diagnostic questions.
@@ 4.1.2. Role-Playing
For the Simple Role-Playing task, we constructed the dataset by adapting the methodology from ALOE  [^5] https://github.com/ShujinWu-0814/ALOE, focusing on user profile and target personality. The training set includes annotations generated by the ChatGPT-4o model (Hurst2024GPT4oSC). The evaluation set, a subset of this dataset, contains conversation contexts grounded in user profile descriptions and specific persona. Generated responses were evaluated based on their ability to effectively simulate the described profile and persona within these contexts.
@@
We adopt the same evaluation methodology as for the MBTI personality simulation task, using LLM-as-a-Judge (Zheng2023JudgingLW). The evaluation dimensions are as follows:
@@ 4.2. Baselines
Preference Prompting (PP) (Jang2023PersonalizedSP) implements personality conditioning through explicit attribute descriptions in prompt engineering, applied across base model, single-preference model, and specialized model(use in Sec. 4.6) configurations. DPO Soups (a variant of Personalized Soups) (Jang2023PersonalizedSP; Zhou2023BeyondOA) acquires the LoRA parameters of independently - trained single - preference DPO models. Following weighted averaging, these parameters are applied to $\pi_{\text{ref}}$ to generate text with multi-personality. MOD (Shi2024DecodingTimeLM) performs a weighted linear combination of the output logits from single - attribute DPO models during decoding and samples based on the combined logits. SCR: Our Speculative Chunk-level based Rejection sampling with parallel multi-preference scoring.
@@ 4.6. Leveraging Specialized Reference Models
To investigate this, we evaluated the performance of SCR when using a different, more specialized reference model. Specifically, we replaced the default base model with a strong, open-source model pre-trained for Role-Playing (Cui2023MachineMA; Wang2025CoSERCL) [^6] https://modelscope.cn/organization/FarReelAILab, [^7] https://huggingface.co/Neph0s/CoSER-Llama-3.1-8B, while keeping our DPO-trained preference models $\pi_{d_{i}}$ unchanged.